OpenStudy (anonymous):

If f(x) = 4x – 1 and g(x) = 3x + 4, what is f(g(x)) ?

OpenStudy (anonymous):

f(g(x)) = 4(3x+4) - 1 f(g(x)) = 12x + 16 - 1 f(g(x)) = 12x + 15